Table 7-6                      Output per hour              Production and   &n

Table 7-6

                     Output per hour              Production and                  Production

                           of work           Consumption without Trade         with Trade

Swords

Belts

Swords

Belts

Swords

Belts

Estonia

5

3

100

40

200

0

Morocco

2

2

60

60

0

120

Estonia and Morocco can produce both swords and belts. Each country has a total of 40 available labor hours for the production of swords and belts. Table 7-6 shows the output per hour of work, the production and consumption quantities without trade, and the production numbers with trade

33) Refer to Table 7-6. With trade, what is the total gain in belt production?

A) 0

B) 20

C) 40

D) 120

34) Refer to Table 7-6. What is the opportunity cost to produce 1 belt in Estonia?

A) 1/3 of a sword

B) 3/5 of a sword

C) 1.67 swords

D) 5 swords
Answer #1

With trade, total number of belts produced =0+120=120.

Without trade the number of belts produced would have been 40+60=100.

Gain =120-100=20.

Opportunity cost to produce one belt in Estonia = 5/3 swords =1.67 swords.

This is because in an hour, Estonia can produce either 3 belts or 5 swords. Hence, for 3 belts 5 swords have to be given up. Similarly for 1 belt, 5/3 swords have to be given up.